OpenStudy (mckenzieandjesus):

A ball is thrown into the air with an upward velocity of 36 ft/s. Its height h in feet after t seconds is given by the function mc016-1.jpg. In how many seconds does the ball reach its maximum height? Round to the nearest hundredth if necessary. What is the ball’s maximum height? 2.25 s; 10 ft 1.13 s; 32.5 ft 1.13 s; 30.25 ft 1.13 s; 70.75 ft

OpenStudy (mckenzieandjesus):

the function is h=-16t^2+36t+10

OpenStudy (mckenzieandjesus):

@Mehek14

Mehek (mehek14):

use the formula \(\sf{\dfrac{-b}{2a}}\)

OpenStudy (mckenzieandjesus):

how do i know which numbers to put into the formula?

Mehek (mehek14):

\(\sf{ax^2+bx+c}\)

Mehek (mehek14):

so for a, you have -16 for b, you have 36

Mehek (mehek14):

\(\dfrac{-36}{2*-16}\)

OpenStudy (mckenzieandjesus):

i got 288?

OpenStudy (mckenzieandjesus):

Ok but those arent on my answer choices

Mehek (mehek14):

it's actually 1.13

OpenStudy (mckenzieandjesus):

would it be 30.25?

Mehek (mehek14):

yes

OpenStudy (mckenzieandjesus):

ok thank you

Mehek (mehek14):

yw

OpenStudy (mckenzieandjesus):

i can't figure out a few other questions.. can u help me?

Mehek (mehek14):

yea

OpenStudy (mckenzieandjesus):

Which of the quadratic functions has the narrowest graph? y=1/3 x^2 y=1/8 x^2 y=-2x^2 y=-3x^2

Mehek (mehek14):

the last one

OpenStudy (mckenzieandjesus):

Ok that is what i was thinking

OpenStudy (mckenzieandjesus):

Which of the quadratic functions has the narrowest graph? 12.6 in. 6.3 in. 10.8 in. 8.9 in.

Mehek (mehek14):

Which of the quadratic functions has the narrowest graph?? I think that's for the other question

OpenStudy (mckenzieandjesus):

Find the value of x. If necessary, round to the nearest tenth.

OpenStudy (mckenzieandjesus):

oops sorry

OpenStudy (mckenzieandjesus):

dont i just times?

Mehek (mehek14):

8.9 is the closest it would be 39.6 which is the closest to 40

OpenStudy (mckenzieandjesus):

are u sure it isnt 6.3 in?

OpenStudy (mckenzieandjesus):

did u times or how did u figure it out?

Mehek (mehek14):

6.3*6.3 = 39.69 39.69/2=19.845

Mehek (mehek14):

formula\(\dfrac{base*height}{2}\)

OpenStudy (anonymous):

mehek u got dis?

OpenStudy (anonymous):

@Mehek14

OpenStudy (mckenzieandjesus):

so 8.9*8.9=79.21 79.21/2=39.605

OpenStudy (mckenzieandjesus):

So am i right or is it 8.9?

Mehek (mehek14):

yes 8.9

OpenStudy (mckenzieandjesus):

OKay thank you!! I dont understand how to do this Write the equation in standard form. Then factor the left side of the equation. 2x2 = 28 – x

OpenStudy (mckenzieandjesus):

wouldn't it be A? (2x + 7)(x – 4) = 0 (2x + 4)(x + 7) = 0 (2x – 7)(x + 4) = 0 (2x – 4)(x + 7) = 0

Mehek (mehek14):

C

Mehek (mehek14):

because the solutions are x = -4 x = 7/2 = 3.500

OpenStudy (mckenzieandjesus):

ohh ok

OpenStudy (mckenzieandjesus):

I have one similar to my very fist question.. A ball is thrown into the air with an upward velocity of 48 ft/s. Its height h in feet after t seconds is given by the function h=-16t^2+48t+6 In how many seconds does the ball reach its maximum height? Round to the nearest hundredth if necessary. What is the ball’s maximum height?

OpenStudy (mckenzieandjesus):

would i do \[48/2*-16?\]

Mehek (mehek14):

-48/2*-16

OpenStudy (mckenzieandjesus):

i get 384

Mehek (mehek14):

-48/-32=1.5

OpenStudy (mckenzieandjesus):

what is the ft?

OpenStudy (mckenzieandjesus):

42 or 54?

Mehek (mehek14):

1.5 is the time

OpenStudy (mckenzieandjesus):

Yeah I know, I just dont know the feet

OpenStudy (mckenzieandjesus):

actually it could be 114 too. 1.5 s; 42 ft 3 s; 6 ft 1.5 s; 54 ft 1.5 s; 114 ft

Mehek (mehek14):

plug 1.5 for t in the equation

OpenStudy (mckenzieandjesus):

114 :)

Mehek (mehek14):

so D

OpenStudy (mckenzieandjesus):

i did 4*pi*24^2 and got 7238.229474 is that right? A sports recreation company plans to manufacture a beach ball with a surface area of 7238 in.2 Find the radius of the beach ball. Use the formula A=4 pi r^2 where A is the surface area and r is the radius of the sphere. 576 in. 75 in. 24 in. 48 in